Find the​ (a) mean,​ (b) median,​ (c) mode, and​ (d) midrange for the given sample data.

An experiment was conducted to determine whether a deficiency of carbon dioxide in the soil affects the phenotype of peas. Listed below are the phenotype codes where

1 equals smooth dash yellow1=smooth-yellow​,

2 equals smooth dash green2=smooth-green​,

3 equals wrinkled dash yellow3=wrinkled-yellow​,

and

4 equals wrinkled dash green4=wrinkled-green.

Do the results make​ sense?

44

11

44

22

44

44

11

44

11

44

22

11

22

33

​(a) The mean phenotype code is

nothing.

​(Round to the nearest tenth as​ needed.)

​(b) The median phenotype code is

nothing.

​(Type an integer or a​ decimal.)

​(c) Select the correct choice below and fill in any answer boxes within your choice.

A.The mode phenotype code is

nothing.

​(Use a comma to separate answers as​ needed.)

B.

There is no mode.

​(d) The midrange of the phenotype codes is

nothing.

​(Type an integer or a​ decimal.)

Do the measures of center make​ sense?

A.

Only the​ mean, median, and mode make sense since the data is numerical.

B.

Only the mode makes sense since the data is nominal.

C.

Only the​ mean, median, and midrange make sense since the data is nominal.

D.

All the measures of center make sense since the data is numerical.
Answer #1

The statistical software output for this problem is:

Summary statistics: Column Mean Median Mode Min Max Data 2.6428571 2.5 4 1 4

Hence,

a) Mean = 2.6

b) Median = 2.5

c) Mode = 4

d) Midrange = (Min + Max)/2 = (1 + 4)/2 = 2.5

Only the mode makes sense since the data is nominal. Option B is correct.